# A distribution-guided Mapper algorithm

### v2

@@ 3 Results
In this section, we compare our proposed D-Mapper algorithm with the classic Mapper algorithm via several experiments. In this paper, we implement D-Mapper by expanding the Mapper algorithm in the KeplerMapper package version 2.0.1 [21] in Python version 3.11.0. The extended persistence diagram and bottleneck bootstrap are computed by a Python package GUDHI version 3.8.0 [22].
We compare our proposed D-Mapper algorithm with the classic Mapper algorithm using the metric $SC_{adj}$ and $SC_{norm}$. We set the interval number $n$ to be identical for both algorithms and use the same clustering algorithm within each hypercube. The clustering method is the density-based spatial clustering of applications with noise (DBSCAN) implemented in the scikit-learn library version 1.1.3. We use a fixed grid to find the best model concerning the $SC_{adj}$ for both the D-Mapper and the classic Mapper. $50$ equally spaced grids in range $(0,\alpha^{\prime})$ are used to select parameter $\alpha$ in the D-Mapper. Similarly, $50$ equally spaced grids in range $(0,0.5)$ are used to tune the overlapping percentage $p$ in the classic Mapper algorithm. The bottleneck bootstrap sampling steps are set to $100$ to compute $85$ confidence level. In this section, we compare the D-Mapper with the classic Mapper via several examples.
@@ 3.1 Two disjoint circles
This dataset is created by sampling points randomly from two disjoint circles with centers $(0,0)$ and $(3,0)$ and a radius of $1$. $5000$ points are uniformly sampled on each circle. Figure 6 (a) provides a visualization of the sampled data points. This dataset has a distinctive shape and provides a straightforward performance comparison between the classical Mapper and D-Mapper algorithms.
In this experiment, we choose the filter function to be a function that projects the original data onto the $X$-axis. We set the number of intervals $n$ to $12$. The clustering algorithm is DBSCAN with a radius of $0.5$ and a minimum of samples $3$. The parameters $p$ of the classic Mapper algorithm and parameter $\alpha$ of the D-Mapper are tuned via grid search, $p=0.02$ and $\alpha=0.127$. The comparison of different evaluation metrics is shown in Table 1 and Figure 4. In Figure 4, the color of a node indicates the average value of the projected data within the node. The output graphs of the classic Mapper and D-Mapper are similar, and both algorithms capture the topological features of the dataset effectively without noise, result in a $TSR$ of 1 for both algorithms. The D-Mapper performs better than the classic Mapper in terms of clustering, as evidenced by its higher $SC_{norm}$. This indicates that D-Mapper has an advantage in identifying more meaningful clusters. Moreover, we show cases that both Mapper algorithms concerned on $SC$ only in Figure 4 (c) and (d). In these two cases the $SC_{norm}$ are higher but the $TSR$ are lower than the results concerned on $SC_{adj}$. The topology structures of these two cases are obviously different from the dataset in these two figures. The distinct results in this example validate the utility of our proposed metric. It provides a quantitative approach for evaluating Mapper-type algorithms from both topological signal preservation and clustering.
Table 1: Results of the D-Mapper and classic Mapper on two disjoint circles dataset.The $1$st and $2$nd rows are results that with the larger $SC_{adj}$. The $3$rd and $4$th rows are cases when the output graphs have larger $SC$ values, but lower $TSR$s.
@@ 3.2 Two intersecting circles
In this section, we compare the performance of the D-Mapper and the classic Mapper on a two intersecting circles dataset shown in Figure 6 (b). The data points are generated from two intersecting circles with a radius of $1$ and centers $(0,0)$ and $(1.5,0)$, respectively. The data generating process is similar to the previous example, $5000$ points are sampled from each circle. The filter function and clustering algorithm is the same as in the previous example. The number of intervals $n$ is set to $8$. The parameters $p$ of the classic Mapper algorithm and parameter $\alpha$ of the D-Mapper are tuned to $p=0.02$ and $\alpha=0.088$. The results are given in Figure 5 and Table 2. As shown in Figure 5 (a,b), the D-Mapper outperforms the classic Mapper concerned on $SC$ scores (or the $SC_{adj}$) and has the same topological signal rate as the classic Mapper. Similar to the two disjoint circles example, Figure 5 (c) and (d) also give cases when the metric $SC$ fails to work.
Table 2: Results of the D-Mapper and classic Mapper on two intersecting circles dataset. The $1$st and $2$nd rows are results that with the larger $SC_{adj}$. The $3$rd and $4$th rows are cases when the output graphs have larger $SC$ values, but lower $TSR$s.
@@
Figure 6: (a) Upper panel: visualization of two disjointed unit circles whose centers are (0,0) and (3,0) respectively. Bottom panel: the intervals produced by both Mappers. The intervals produced by D-Mapper have a gap between two circles, these intervals are more reasonable than classic Mapper’s. Because there are no points between two disjointed circles, the intervals between two circles are meaningless. (b) Upper panel: a visualization of two intersecting unit circles dataset whose centers are $(0,0)$ and $(1.5,0)$ respectively. Bottom panel: the intervals produced by both Mappers. The intervals of D-Mapper can assign more reasonable intervals than classic Mapper on the intersecting part.
@@
In this experiment, we compare the D-Mapper algorithm with the classic Mapper algorithm on a 3D cat dataset with a more complex topology structure. The 3D cat dataset is originally created by [23]. Figure 7 (a) shows a visualization of the 3D cat dataset. The filter function is the sum of all coordinates. The parameters of the classic Mapper are set to $n=9,~{}p=0.32$; and the parameters of the D-Mapper are set to $n=9,~{}\alpha=0.01$. The clustering method is DBSCAN with a radius of $0.1$ and a minimum of samples $5$. The results and evaluation metrics are shown in Figure 7 and Table 3.
The D-Mapper algorithm results higher $SC_{norm}$ scores and the same $TSR$ compares to the classic Mapper algorithm. The output graph of the D-Mapper has one more loop than the classic Mapper, which represents the relationship between the ears, face (including jaw) and neck of the cat (yellow, olive green and light green dots in Figure 7 c-d). This example also shows that the D-Mapper could capture subtle features of complex objects better than the classic Mapper due to the flexible intervals and overlap ratios.
@@
Figure 7: Results of the classic Mapper and D-Mapper on the 3D cat dataset. (a) Visualization of the 3D cat dataset. (b) The intervals produced by both Mappers. The intervals of both Mappers are very close except intervals near to $1$, this difference gives more details on cat head in D-Mapper. (c) The graph of the classic Mapper with the largest $SC_{adj}(0.740)$: $n=9,p=0.32$. (d) The graph of the D-Mapper with the largest $SC_{adj}(0.755)$: $n=9,\alpha=0.01$.
@@ 3.4 Covid-19 dataset
In this section, we apply our proposed method to a real dataset. The Covid-19 pandemic poses an enormous threat to global health and economics. To combat the epidemic, one of the fundamental tasks is identifying and monitoring the mutation of the SARS-COV-2 coronavirus. Studying virus evolution helps understand the basic biological characteristics, develop vaccines and drugs, and forecast trends. A traditional way to study viral mutations is to construct phylogenetic trees based on genetic sequences. However, tree structure representations focus on vertical processes. Whereas, some processes are horizontal. For viruses, homologous recombination and reassortment are typical ways that can lead to non-tree-like representation [24]. We employed the D-Mapper algorithm to analyze a dataset of SARS-CoV-2 coronavirus RNA sequences, utilizing it as a rapid exploratory data analysis tool. This approach allowed us to examine viral mutations’ vertical and horizontal evolutionary processes across the entire genomic scale. The dataset contains $357$ SARS-COV-2 coronavirus whole RNA sequences with different lineages from GenBase (https://ngdc.cncb.ac.cn/genbase/) in National Genomics Data Center [25].
@@
The number of intervals are set to $n=15,\alpha=0.006$ via grid tuning in the D-Mapper algorithm. The DBSCAN with a radius of $0.6$ and a minimum of samples $3$ is chosen in the clustering algorithm. The results are shown in Figure 8, and the full list is provided in the supplementary. The output structure represents the evolution process of these lineages. The loop structure represents the horizontal evolution process. The lineages of the two isolated nodes may indicate significant differences from others, these lineages may warrant further investigation. The result of classic Mapper algorithm is also provided in the appendix.
Figure 8: The result of D-Mapper on SARS-COV-2 dataset. There are $19$ nodes in total, indexed from 00 to $18$.
Some important lineages in these nodes are shown in Table 4. The bottom green loop nodes (node 9, 10, 12) contain many XBB variants, which form the recombination of two subvariants and cause large-scale infections worldwide [30]. Nodes 4 and 5 contain the JN.1 variant, which is currently widespread worldwide [31]. When compared to its predecessor, BA.2.86, the JN.1 variant only has one additional mutation within its spike protein. Nonetheless, BA.2.86 descent while JN.1 rapidly become predominant [32]. The D-Mapper graph can give some insights into this phenomenon, where BA.2.86 and JN.1 are both in a loop structure (nodes 5, 6, 8), which indicates the existence of complex horizontal evolution between these two variants. Moreover, JN.1 variant is located in both linear (node 4) and cyclical structures (node 5), suggesting it may play a pivotal role in the evolutionary process. Consequently, other variants presented at these nodes (nodes 4, 5) warrant further scrutiny. The upper purple loop (nodes 0, 1, 2, 3) encompasses variant BA.2, which is the predecessor of BA.2.86. The spike protein of BA.2.86 has more than 30 mutations than BA.2 [33], which are reflected as significant distances on the D-Mapper graph. In summary, the D-Mapper algorithm offers crucial insights into the rapid and complex evolution process of viruses.
@@ 4 Discussion
Our proposed algorithm is a probabilistic model-based approach that uses data intrinsic characteristics and probability models to generate distribution-guided covers and improved topological features. It is a viable alternative to non-probabilistic approaches. With the D-Mapper, once we get the distribution of the projected data, we further can get the intervals from each distribution. Theoretically, the distribution of original data can be computed by transformation of distributions if the filter function is well proposed, this may be helpful for further theoretical analysis.
@@ 5 Conclusion
In this paper, we propose a distribution-guided Mapper algorithm (D-Mapper) to relax the fixed intervals and overlapping ratios restriction in the classic Mapper algorithm. Our proposed algorithm combines a mixture model and the Mapper algorithm to obtain irregular intervals based on the density of the projected data. With these irregular intervals, the D-Mapper algorithm can gain deeper topological insights and enhance clustering outcomes. To validate the effectiveness of our proposed algorithm, we introduce the $SC_{adj}$ score to combine $SC$ scores and extended persistence diagram as a metric to reflect the performance of overlap clustering and persistent homology.
We also conduct simulation studies with different complexity to evaluate the D-Mapper algorithm by the $SC_{adj}$. In all simulations, the D-Mapper outperforms the classic Mapper algorithm concerning on the metric $SC_{adj}$. The $SC_{norm}$ of the D-Mapper are all higher than the classic Mapper algorithm, and all $TSR$s are $1$. This indicates that the D-Mapper can achieve better clustering than the classic Mapper algorithm while outputting high-quality Reeb graph approximations. Note that we tune parameters by grid searching the highest $SC_{adj}$, thus the $TSR$ is relatively high. In many cases, the $TSR$ could be low, as we show in Table 1 and Table 2. Our experimental results also indicate that our proposed metric $SC_{adj}$ is more stable than the metric $SC$. We also apply the D-Mapper algorithm to the SARS-COV-2 coronavirus RNA sequences, and the result shows that the D-Mapper algorithm can reflect both the vertical and horizontal evolution processes.
[…]
Acknowledgements This project was supported by the Shanghai Science and Technology Program (No. 21010502500), the startup fund of ShanghaiTech University.
Appendix A Output of the classic Mapper algorithm on Covid-19 dataset
The result of the classic Mapper algorithm is shown in Figure 9, and the full list is provided in the supplementary.
Figure 9: The result of classic Mapper on SARS-COV-2 dataset ($n=15,p=0.49$). There are $20$ nodes in total, indexed from 00 to $19$.
Appendix B Bottleneck bootstrap algorithm
@@
Let the dataset with filtered values be a set of pairs $(X_{1},Y_{1}),...,(X_{n},Y_{n})$. The Mapper graph is represented by $M$ and the bootstrap samples are denoted as $(X^{*}_{1},Y^{*}_{1}),...,(X^{*}_{n},Y^{*}_{n})$. The Mapper graph computed from these bootstrap samples is denoted by $M^{*}$. Given a confidence level $\epsilon$, this algorithm outputs a distance $d_{\epsilon}$ which can be utilized to differentiate between noise and signal in the extended persistence diagram.
Algorithm 4 Bottleneck bootstrap
1:Input: Dataset and filtered value $(X_{1},Y_{1}),...,(X_{n},Y_{n})$, all parameters for the Mapper or D-Mapper algorithm, confidence level $\epsilon$.
2:Computing Mapper graph $M$ by the Mapper or D-Mapper algorithm by using original samples.
@@
5:     Computing bootstrap Mapper graph $M^{*}$ by the Mapper or D-Mapper algorithm.
@@
8:Find $\alpha$ quantile from $\{d_{1},...,d_{N}\}$, denote as $d_{\epsilon}$.

### v3

@@ 3 Results
In this section, we compare our proposed D-Mapper algorithm with the classic Mapper algorithm via several experiments. In this work, we implement D-Mapper by expanding the Mapper algorithm in the KeplerMapper package version 2.0.1 [21] in Python version 3.11.0. The extended persistence diagram and bottleneck bootstrap are computed by a Python package GUDHI version 3.8.0 [22].
We compare our proposed D-Mapper algorithm with the classic Mapper algorithm using the metrics $SC_{adj}$ and $SC_{norm}$. We set the interval number $n$ to be identical for both algorithms and use the same clustering algorithm within each hypercube. The clustering method is the density-based spatial clustering of applications with noise (DBSCAN) implemented in the scikit-learn library version 1.1.3. We use a fixed grid to find the best model concerning the $SC_{adj}$ for both the D-Mapper and the classic Mapper. $50$ equally spaced grids in range $(0,\alpha^{\prime})$ are used to select parameter $\alpha$ in the D-Mapper. Similarly, $50$ equally spaced grids in range $(0,0.5)$ are used to tune the overlapping percentage $p$ in the classic Mapper algorithm. The bottleneck bootstrap sampling steps are set to $100$ to compute the $85\%$ quantile. In this section, we compare the D-Mapper with the classic Mapper via several examples.
@@ 3.1 Two disjoint circles
This dataset is created by sampling points randomly from two disjoint circles with centers $(0,0)$ and $(3,0)$ and a radius of $1$. $5000$ points are uniformly sampled on each circle. The upper panel of Figure 6 (a) provides a visualization of the sampled data points. This dataset has a distinctive shape and provides a straightforward performance comparison between the classic Mapper and D-Mapper algorithms.
In this experiment, we choose the filter function to be a function that projects the original data onto the $X$-axis. We set the number of intervals $n$ to $12$. The clustering algorithm is DBSCAN with a radius of $0.5$ and a minimum of samples $3$. The parameters $p$ of the classic Mapper algorithm and parameter $\alpha$ of the D-Mapper are tuned via grid search, $p=0.02$ and $\alpha=0.127$. The comparison of different evaluation metrics is shown in Table 1 and Figure 4. The resulting intervals are given in the lower panel of Figure 6 (a). In Figure 4, the color of a node indicates the average value of the projected data within the node. The output graphs of the classic Mapper and D-Mapper are similar (Figure 4 a-b), and both algorithms capture the topological features of the dataset effectively without noise, resulting in a $TSR$ of 1 for both algorithms. The D-Mapper performs better than the classic Mapper in terms of clustering, as evidenced by its higher $SC_{norm}$. This indicates that D-Mapper has an advantage in identifying more meaningful clusters. Moreover, we display cases in both the classic Mapper and D-Mapper algorithms concerned with the $SC$ only in Figure 4 (c) and (d). In these two cases, the $SC_{norm}$ values are higher but the $TSR$ values are lower than the results concerned with the $SC_{adj}$. The topology structures of these two cases are different from the dataset in these two figures. The distinct results in this example validate the utility of our proposed metric. It provides a quantitative approach for evaluating Mapper-type algorithms from both topological signal preservation and clustering.
Table 1: Results of the D-Mapper and classic Mapper on the two disjoint circles dataset. The $1$st and $2$nd rows are results with the larger $SC_{adj}$. The $3$rd and $4$th rows are cases when the output graphs have larger $SC$ values but lower $TSR$s.
@@ 3.2 Two intersecting circles
In this section, we compare the performance of the D-Mapper and the classic Mapper on a two intersecting circles dataset shown in the upper panel of Figure 6 (b). The data points are generated from two intersecting circles with a radius of $1$ and centers $(0,0)$ and $(1.5,0)$, respectively. The data generating process is similar to the previous example, $5000$ points are sampled from each circle. The filter function and clustering algorithm are the same as in the previous example. The number of intervals $n$ is set to $8$. The parameters $p$ of the classic Mapper algorithm and parameter $\alpha$ of the D-Mapper are tuned to $p=0.02$ and $\alpha=0.088$. The results are given in Figure 5 and Table 2. The resulting intervals are given in the lower panel of Figure 6 (b). As shown in Figure 5 (a,b), the D-Mapper outperforms the classic Mapper concerning the $SC$ scores (or the $SC_{adj}$) and has the same topological signal rate as the classic Mapper. Similar to the two disjoint circles example, Figure 5 (c) and (d) also give cases when the metric $SC$ fails to work.
Table 2: Results of the D-Mapper and classic Mapper on the two intersecting circles dataset. The $1$st and $2$nd rows are results with the larger $SC_{adj}$. The $3$rd and $4$th rows are cases when the output graphs have larger $SC$ values but lower $TSR$s.
@@
Figure 6: (a) Upper panel: visualization of two disjointed unit circles whose centers are (0,0) and (3,0), respectively. Bottom panel: the intervals produced by both the classic Mapper and D-Mapper. The intervals produced by the D-Mapper have a gap between two circles, these intervals are more reasonable than the classic Mapper’s. As there are no points between the two disjointed circles, the intervals between the two circles are meaningless. (b) Upper panel: a visualization of two intersecting unit circles dataset whose centers are $(0,0)$ and $(1.5,0)$, respectively. Bottom panel: the intervals produced by both the classic Mapper and D-Mapper. The intervals of the D-Mapper can assign more reasonable intervals than the classic Mapper on the intersecting part.
@@
In this experiment, we compare the D-Mapper algorithm with the classic Mapper algorithm on a 3D cat dataset with a more complex topology structure. The 3D cat dataset is originally created by [23]. Figure 7 (a) shows a visualization of the 3D cat dataset. The filter function is the sum of all coordinates. The parameters of the classic Mapper are set to $n=9,~{}p=0.32$; and the parameters of the D-Mapper are set to $n=9,~{}\alpha=0.01$. The clustering method is DBSCAN with a radius of $0.1$ and a minimum of samples $5$. The results and evaluation metrics are shown in Figure 7 (b,c,d) and Table 3.
The D-Mapper algorithm results in higher $SC_{norm}$ scores and the same $TSR$ compared to the classic Mapper algorithm. The output graph of the D-Mapper has one more loop than the classic Mapper, which represents the relationship between the ears, face (including jaw) and neck of the cat (yellow, olive green and light green dots in Figure 7 c-d). This example also shows that the D-Mapper could capture subtle features of complex objects better than the classic Mapper due to the flexible intervals and overlapping ratios.
@@
Figure 7: Results of the classic Mapper and D-Mapper on the 3D cat dataset. (a) Visualization of the 3D cat dataset. (b) The intervals produced by both the classic Mapper and D-Mapper. The intervals of both the classic Mapper and D-Mapper are very similar except for intervals near $1$; this difference gives more details on the cat head in D-Mapper. (c) The graph of the classic Mapper with the largest $SC_{adj}(0.740)$: $n=9,p=0.32$. (d) The graph of the D-Mapper with the largest $SC_{adj}(0.755)$: $n=9,\alpha=0.01$.
@@ 3.4 Covid-19 dataset
In this section, we apply our proposed method to a real dataset. The Covid-19 pandemic poses an enormous threat to global health and economics. To combat the epidemic, one of the fundamental tasks is identifying and monitoring the mutation of the SARS-COV-2 coronavirus. Studying virus evolution helps understand the basic biological characteristics, develop vaccines and drugs, and forecast trends. A traditional way to study viral mutations is to construct phylogenetic trees based on genetic sequences. However, tree structure representations focus on vertical processes. Whereas, some processes are horizontal. For viruses, homologous recombination and reassortment are typical ways that can lead to non-tree-like representation [24]. We employ the D-Mapper algorithm to analyze a dataset of SARS-CoV-2 coronavirus RNA sequences, utilizing it as a rapid exploratory data analysis tool. This approach allows us to examine viral mutations’ vertical and horizontal evolutionary processes across the entire genomic scale. The dataset contains $357$ SARS-COV-2 coronavirus whole RNA sequences with different lineages from GenBase (https://ngdc.cncb.ac.cn/genbase/) in National Genomics Data Center [25].
@@
The number of intervals are set to $n=15,\alpha=0.006$ via grid tuning in the D-Mapper algorithm. The DBSCAN with a radius of $0.6$ and a minimum of samples $3$ is chosen in the clustering algorithm. The results are shown in Figure 8, and the full list is provided in the supplementary file. The output structure represents the evolutionary process of these lineages. The loop structure represents the horizontal evolutionary process. The lineages of the two isolated nodes may indicate significant differences from others, these lineages may warrant further investigation. In addition, the result of the classic Mapper algorithm is provided in Figure 9 in Appendix B.
Figure 8: The result of D-Mapper on the SARS-COV-2 dataset. There are $19$ nodes in total, indexed from 00 to $18$.
Some important lineages in these nodes are shown in Table 4. The bottom green loop nodes (nodes 9, 10, 12) contain many XBB variants, which form the recombination of two subvariants and cause large-scale infections worldwide [30]. Nodes 4 and 5 contain the JN.1 variant, which is currently widespread worldwide [31]. When compared to its predecessor, BA.2.86, the JN.1 variant only has one additional mutation within its spike protein. Nonetheless, BA.2.86 descent while JN.1 rapidly become predominant [32]. The D-Mapper graph can give some insights into this phenomenon, where BA.2.86 and JN.1 are both in a loop structure (nodes 5, 6, 8), which indicates the existence of complex horizontal evolution between these two variants. Moreover, JN.1 variant is located in both linear (node 4) and cyclical structures (node 5), suggesting it may play a pivotal role in the evolutionary process. Consequently, other variants presented at these nodes (nodes 4, 5) warrant further scrutiny. The upper purple loop (nodes 0, 1, 2, 3) encompasses variant BA.2, which is the predecessor of BA.2.86. The spike protein of BA.2.86 has more than 30 mutations than BA.2 [33], which are reflected as significant distances on the D-Mapper graph. In summary, the D-Mapper algorithm offers crucial insights into the rapid and complex evolutionary process of viruses.
@@ 4 Discussion
Our proposed algorithm is a probabilistic model-based approach that uses data intrinsic characteristics and probability models to generate distribution-guided covers and improved topological features. It is a viable alternative to non-probabilistic approaches. With the D-Mapper, once we get the distribution of the projected data, we can further get the intervals from each distribution. Theoretically, the distribution of original data can be computed by transformation of distributions if the filter function is well proposed, this may be helpful for further theoretical analysis.
@@ 5 Conclusion
In this work, we propose a distribution-guided Mapper algorithm (D-Mapper) to relax the fixed intervals and overlapping ratios restriction in the classic Mapper algorithm. Our proposed algorithm combines a mixture model and the Mapper algorithm to obtain irregular intervals based on the density of the projected data. With these irregular intervals, the D-Mapper algorithm can gain deeper topological insights and enhance clustering outcomes. To validate the effectiveness of our proposed algorithm, we introduce the $SC_{adj}$ score to combine $SC$ scores and extended persistence diagram as a metric to reflect the performance of overlap clustering and persistent homology.
We also conduct simulation studies with different complexity to evaluate the D-Mapper algorithm by the $SC_{adj}$. In all simulations, the D-Mapper outperforms the classic Mapper algorithm concerning the metric $SC_{adj}$. The $SC_{norm}$ of the D-Mapper are all higher than the classic Mapper algorithm, and all $TSR$s are $1$. This indicates that the D-Mapper can achieve better clustering than the classic Mapper algorithm while outputting high-quality Reeb graph approximations. Note that we tune parameters by grid searching the highest $SC_{adj}$, thus the $TSR$ is relatively high. In many cases, the $TSR$ could be low, as we show in Table 1 and Table 2. Our experimental results also indicate that our proposed metric $SC_{adj}$ is more stable than the metric $SC$. We also apply the D-Mapper algorithm to the SARS-COV-2 coronavirus RNA sequences, and the result shows that the D-Mapper algorithm can reflect both the vertical and horizontal evolutionary processes.
Appendix A The bottleneck bootstrap algorithm
@@
Paired the original data points and the filtered values as $(X_{1},Y_{1}),...,(X_{n},Y_{n})$. The Mapper graph is represented by $M$ and the bootstrap samples are denoted as $(X^{*}_{1},Y^{*}_{1}),...,(X^{*}_{n},Y^{*}_{n})$. The Mapper graph generated from these bootstrap samples is denoted by $M^{*}$. Given a confidence level $\epsilon$, this algorithm outputs a distance $d_{\epsilon}$, which can be utilized to differentiate between noise and signal in the extended persistence diagram. Algorithm 4 gives the details of the bottleneck bootstrap algorithm.
Algorithm 4 The bottleneck bootstrap algorihtm
1:Input: Paired orignial data and filtered values $(X_{1},Y_{1}),...,(X_{n},Y_{n})$, all parameters for the Mapper or D-Mapper algorithm, a confidence level $\epsilon$.
2:Generating a Mapper graph $M$ by the Mapper or D-Mapper algorithm by using original samples.
@@
5:     Generating a bootstrap Mapper graph $M^{*}$ by the Mapper or D-Mapper algorithm.
@@
8:Find the $\epsilon$ quantile from $\{d_{1},...,d_{N}\}$, denote as $d_{\epsilon}$.
@@
Appendix B The output of the classic Mapper algorithm on the Covid-19 dataset
The result of the classic Mapper algorithm is shown in Figure 9, and the full list is provided in the supplementary file.
Figure 9: The result of the classic Mapper on the SARS-COV-2 dataset ($n=15,p=0.49$). There are $20$ nodes in total, indexed from 00 to $19$.
Acknowledgements
The authors thank ShanghaiTech University for supporting this work through the startup fund and the HPC Platform.
Authors’ contributions
YT—conceptualization, data curation, formal analysis, methodology, coding, visualization, writing—original draft. SG—conceptualization, data curation, formal analysis, methodology, coding evaluation, project administration, supervision, writing—final draft, writing—review and editing. All authors read and approved the final manuscript.
Funding
YT and SG are supported by the Shanghai Science and Technology Program (No. 21010502500), the National Natural Science Foundation of China (12401383).
Availability of data and materials
The paper employs publicly accessible datasets in the 3D cat and Covid-19 experiments. The simulated data of the two disjoint circles and two intersecting circles, the results of the all experiments and code are available in the public repository: https://github.com/ShufeiGe/D-Mapper.
Declarations
Ethical approval and consent to participate
Not applicable.
Consent for publication
Not applicable.
Conflict of Interests
The authors declare that they have no competing interests.